OpenStudy (anonymous):

suppose that the x-intercept of the graph of y=f(x) are -8 and 1. what are the x-intercept of the graph of y=f(x+4)

OpenStudy (dangerousjesse):

\(y = f(x) = (x+8)(x-1)=x^2+7 x-8\)

OpenStudy (dangerousjesse):

Now solve \(x^2+7x-8\) using the quadratic equation.

OpenStudy (loser66):

f(x) = (x+8)(x-1) f(x+4) = ((x+4)+8)((x+4)-1) Now, solve for x :)